# Resolving the Orientations of and Separation between an Overlapping Pair of Dipole Emitters

### v1

@@ Abstract
We prove that it is impossible to distinguish between two fluorescent dipoles versus a single rotating molecule by either modulating the polarization of pumping light or the detection dipole-spread function (DSF). Even if the target is known to be a dipole pair, existing orientation-imaging methods perform poorly for measuring their angular separation. We propose modulating both excitation polarization and the DSF simultaneously, thereby demonstrating robust discrimination between dipole pairs versus single molecules. Combined excitation modulation-DSF engineering improves the precision of measuring centroid orientation by 50% and angular separation by 4- to 6-fold.
@@
Single-molecule (SM) nanoscopy has become invaluable for overcoming the optical diffraction limit and observing nanoscale structures and dynamics within biological systems [1, 2]. Leveraging the resolvability of blinking molecules over time, these methods repeatedly detect and localize isolated point-spread functions (PSFs) to determine molecular positions [3] and/or orientations [4, 5]. However, leveraging molecular blinking necessarily sacrifices temporal resolution [6] for improved spatial resolution, thereby limiting the ability to observe dynamic processes in biological systems [7, 8]. Resolving emitters beyond the Abbé diffraction limit remains an active research area [9, 10].
@@
Recent studies have established fundamental quantum estimation limits for resolving two incoherent point sources in 2D [11, 12, 13] and 3D [14, 15, 16], and several demonstrations have exhibited resolution performance approaching theoretical limits [17, 18, 19]. As MINFLUX and related techniques [20, 21, 22, 23] approach ångström-level resolution, here, we explore an orthogonal approach: how well can two spatially overlapping dipole emitters be resolved using polarized light? Modulating the polarization of an illumination beam has been demonstrated to improve spatial resolution [24, 25, 26, 27, 28], but to our knowledge, the physical limits of resolving a pair of dipole emitters based upon their separation in orientation space has yet to be established. Given numerous recent developments in measuring dipole orientations [27, 29, 30, 31, 28, 32, 33, 34, 35, 36], such a theory would provide tremendous guidance for improving measurements of single-molecule rotational dynamics at the nanoscale.
@@
In this work, we derive a simple mathematical proof showing that any technique that solely 1) modulates the polarization of pumping light, termed excitation modulation (ExM), or 2) modulates the phase and/or polarization of fluorescence emission, called dipole-spread function (DSF) engineering, cannot distinguish between a pair of dipole emitters versus a single rotating dipole. This degeneracy stems from the fundamental physics of collecting fluorescence photons and cannot be overcome without prior knowledge of the target sample. Further, even if one is certain that an image originates from a pair of dipoles, our analysis shows that all existing techniques cannot measure their separation angle with high precision. To address these issues, we propose combining ExM and DSF engineering to not only improve orientation measurement performance but also overcome the problem of resolving a single wobbling molecule from a pair of overlapping molecules.
@@
Interestingly, we find that if $\phi_{c}=\tilde{\phi}$ and $\cos{\delta}=\operatorname{sinc}{(2\tilde{\omega})}$, then the second moments of the dipole pair and the single wobbling dipole are identical. Thus, any imaging system is fundamentally unable to distinguish between a pair of dipoles from a single dipole, regardless of the centroid orientation $\phi_{c}$. Figure 1(d) shows the one-to-one equivalence between the separation $\delta$ of the dipole pair and wobble angle $\tilde{\omega}$ of the SM. Critically, illuminating the sample with various excitation beam polarizations via ExM also cannot detect any difference between a dipole pair and a single dipole [39]. We further remark that this irresolvability occurs regardless of the precise type of dipole wobble, e.g., uniform rotation, rotation within a harmonic potential, etc. [42, 43, 44].
Figure 2: Precision of measuring the centroid orientation $\left(\theta_{c},\phi_{c}\right)$ of and separation angle $\Delta$ between two fixed dipoles with 100 photons detected. (a) Example dipole pair with $\bm{\mu}_{1}=(\cos{(\phi_{c}-\phi_{\Delta}/2)}\sin{(\theta_{c}-\theta_{\Delta}/2)}$, $\sin{(\phi_{c}-\phi_{\Delta}/2)}\sin{(\theta_{c}-\theta_{\Delta}/2)}$, $\cos{(\theta_{c}-\theta_{\Delta}/2)})$ and $\bm{\mu}_{2}=(\cos{(\phi_{c}+\phi_{\Delta}/2)}\sin{(\theta_{c}+\theta_{\Delta}/2)}$, $\sin{(\phi_{c}+\phi_{\Delta}/2)}\sin{(\theta_{c}+\theta_{\Delta}/2)}$, $\cos{(\theta_{c}+\theta_{\Delta}/2)})$ and separation $\Delta=\arccos{(\bm{\mu}_{1}^{\intercal}\bm{\mu}_{2})}=20^{\circ}$. (b) Best-possible precisions, quantified by the Cramér-Rao bound (CRB) of measuring (i) centroid orientation $\sigma_{c}$ and (ii) separation angle $\sigma_{\Delta}$ for a dipole pair with $\theta_{c}=60^{\circ}$. (c) Best-possible precisions of measuring (i) centroid orientation $\sigma_{c}$ and (ii) separation angle $\sigma_{\Delta}$, averaging over $\phi_{c}\in[0^{\circ},360^{\circ}]$. Blue, direct BFP imaging [40]; yellow, CHIDO [29]; green, polarized vortex DSF [30]; red, pixOL DSF [32]. The gray regions show the precision limit bounded by the quantum CRB (QCRB).
@@
We also use these transformations to compute analogous best-possible precisions for any imaging system using quantum FI (QFI) $\bm{\mathcal{K}}$. Note that while classical FI $\bm{\mathcal{J}}$ is calculated for specific microscopes imaging molecules at specific orientations $\bm{\Theta}$ [49], QFI provides a universal precision bound for any imaging system [46, 13]. While one cannot guarantee that a method can be devised to reach the QFI limit, comparing the CRB achieved by existing techniques to the quantum CRB (QCRB) [39] provides useful context for evaluating performance.
@@
where $a$ incorporates the absorption cross-section and quantum yield of the molecule, wavelength and intensity of the excitation beam, etc. [39, 52] and $\bm{H}\in\mathbb{R}^{NL\times 15}$ represents the imaging system’s response to each fourth moment $q$.
@@
Figure 4: Performance of using combined excitation modulation (ExM) and dipole-spread function (DSF) engineering. (a) Schematic of (i) 3 designed wavevectors $\bm{k}$ and (ii) their associated s- and p-polarized states for ExM. The three $\bm{k}$ vectors lie on a cone oriented along the $\mu_{z}$ direction with half-angle $\beta$, and the relative azimuthal angle between any pair is $120^{\circ}$. (b) Estimation precision for the (i) centroid orientation of and (ii) separation between a pair of dipoles using ExM-pixOL with various $\beta$ angles. Purple, $\beta=30^{\circ}$; orange, $\beta=45^{\circ}$; red, $\beta=\arccos{(1/\sqrt{3})}\approx 54.74^{\circ}$; blue, $\beta=60^{\circ}$. (c) Same as (b) for different DSFs and $\beta=\arccos{(1/\sqrt{3})}\approx 54.74^{\circ}$. Yellow, CHIDO [29]; green, polarized vortex DSF [30]; red, pixOL DSF [32]. The black line in (b) and (c) refers to the QCRB of DSF engineering as presented in Fig. 2. Precisions are calculated for a dipole pair with separation $\Delta=20^{\circ}$ and are averaged over $\phi_{c}\in[0^{\circ},360^{\circ}]$ with 100 photons detected and zero background.
@@
In summary, we analyzed the fundamental constraints for measuring the orientations of a pair of overlapping dipole emitters using polarization-sensitive single-molecule imaging. Since the second moments of a pair of dipole emitters and a rotating molecule are identical, both excitation modulation and DSF engineering techniques cannot distinguish between the two scenarios without prior knowledge. However, by increasing the dimensionality of the measurement to the fourth-moment space via combined ExM and DSF engineering, pairs of dipoles produce camera images that are different from those of a single wobbling molecule, thereby solving the degenerate measurements of existing techniques. We demonstrated that combined ExM-DSF engineering exhibits superior performance for measuring the centroid orientation of and separation between a pair of molecules. Our work lays the foundation for developing optimal techniques for resolving spatially overlapped dipole emitters and precisely measuring their rotational dynamics in biological samples.
@@ Acknowledgements.
We thank Tingting Wu, Weiyan Zhou, and Suraj Deepak Khochare for their helpful discussions. This work was supported by the National Institute of General Medical Sciences of the National Institutes of Health under Award Number R35GM124858.

### v2

@@ Abstract
We prove that it is impossible to distinguish two spatially overlapping fluorescent molecules from a single rotating molecule, even if one modulates the polarization of pumping light or the detection dipole-spread function (DSF). If the target is known to be a dipole pair, existing imaging methods perform poorly for measuring their angular separation. We propose simultaneously modulating the excitation polarization and DSF, which demonstrates robust discrimination between dipole pairs versus single molecules. Our method improves the precision of measuring centroid orientation by 50% and angular separation by 4- to 6-fold over existing techniques.
@@
Single-molecule (SM) nanoscopy has become invaluable for overcoming the optical diffraction limit and observing nanoscale structures and dynamics within biological systems [1, 2]. Leveraging the resolvability of blinking molecules over time, these methods repeatedly detect and localize isolated point-spread functions to determine molecular positions [3] and/or orientations [4, 5]. However, leveraging molecular blinking necessarily sacrifices temporal resolution [6] for improved spatial resolution, thereby limiting the ability to observe dynamic processes in biological systems [7, 8]. Resolving emitters beyond the Abbé diffraction limit remains an active research area [9, 10].
@@
Recent studies have established fundamental quantum estimation limits for resolving two incoherent point sources in 2D [11, 12, 13] and 3D [14, 15, 16], and several demonstrations have exhibited resolution performance approaching theoretical limits [17, 18, 19]. As MINFLUX and related techniques [20, 21, 22, 23] approach ångström-level resolution, here, we explore an orthogonal approach: how well can two spatially overlapping dipole emitters be resolved using polarized light? Modulating the polarization of an illumination beam has been demonstrated to improve spatial resolution [24, 25, 26, 27, 28], but to our knowledge, the physical limits of resolving a pair of dipole emitters based upon their separation in orientation space have yet to be established. Given numerous recent developments in measuring dipole orientations [27, 29, 30, 31, 28, 32, 33, 34, 35, 36], such a theory would provide tremendous guidance for improving measurements of single-molecule rotational dynamics at the nanoscale.
@@
In this work, we derive a simple mathematical proof showing that any technique that solely 1) modulates the polarization of pumping light, termed excitation modulation (ExM), or 2) modulates the phase and/or polarization of fluorescence emission, called dipole-spread function (DSF) engineering, cannot distinguish between a pair of dipole emitters versus a single rotating dipole. This degeneracy stems from the fundamental physics of collecting fluorescence photons and cannot be overcome without prior knowledge of the target sample. Further, even if one is certain that an image originates from a pair of dipoles, our analysis shows that all existing techniques cannot measure their separation angle with high precision. To address these issues, we propose combining ExM and DSF engineering to both overcome the problem of resolving a single wobbling molecule from a pair of overlapping molecules and significantly improve the precision of measuring the centroid orientation of and angular separation between a pair of molecules.
@@
Interestingly, we find that if $\phi_{c}=\tilde{\phi}$ and $\cos{\delta}=\operatorname{sinc}{2\tilde{\omega}}$, then the second moments of the dipole pair and the single wobbling dipole are identical. Thus, any imaging system is fundamentally unable to distinguish between a pair of dipoles from a single dipole, regardless of the centroid orientation $\phi_{c}$. Figure 1(d) shows the one-to-one equivalence between the separation $\delta$ of the dipole pair and wobble angle $\tilde{\omega}$ of the SM. Critically, illuminating the sample with various excitation beam polarizations via ExM also cannot detect any difference between a dipole pair and a single dipole [39]. We further remark that this irresolvability occurs regardless of the precise type of dipole wobble, e.g., uniform rotation, rotation within a harmonic potential, etc. [42, 43, 44].
Figure 2: Precision of measuring the centroid orientation $\left(\theta_{c},\phi_{c}\right)$ of and separation angle $\Delta$ between two fixed dipoles with 100 photons detected. (a) Example dipole pair with $\bm{\mu}_{1}=(\cos{(\phi_{c}-\phi_{\Delta}/2)}\sin{(\theta_{c}-\theta_{\Delta}/2)}$, $\sin{(\phi_{c}-\phi_{\Delta}/2)}\sin{(\theta_{c}-\theta_{\Delta}/2)}$, $\cos{(\theta_{c}-\theta_{\Delta}/2)})$ and $\bm{\mu}_{2}=(\cos{(\phi_{c}+\phi_{\Delta}/2)}\sin{(\theta_{c}+\theta_{\Delta}/2)}$, $\sin{(\phi_{c}+\phi_{\Delta}/2)}\sin{(\theta_{c}+\theta_{\Delta}/2)}$, $\cos{(\theta_{c}+\theta_{\Delta}/2)})$ and separation $\Delta=\arccos{(\bm{\mu}_{1}^{\intercal}\bm{\mu}_{2})}=20^{\circ}$. (b) Best-possible precisions, quantified by the Cramér-Rao bound (CRB) of measuring (i) centroid orientation $\sigma_{c}$ and (ii) separation angle $\sigma_{\Delta}$ for a dipole pair with $\theta_{c}=60^{\circ}$. (c) Best-possible precisions of measuring (i) centroid orientation $\sigma_{c}$ and (ii) separation angle $\sigma_{\Delta}$, averaging over $\phi_{c}\in[0^{\circ},360^{\circ}]$. Red, pixOL DSF [32]; green, polarized vortex DSF [30]; yellow, CHIDO [29]; blue, direct BFP imaging [40]. The gray regions show the precision limit bounded by the quantum CRB (QCRB).
@@
We also use these transformations to compute analogous best-possible precisions for any imaging system using quantum FI (QFI) $\bm{\mathcal{K}}$. Note that while classical FI $\bm{\mathcal{J}}$ is calculated for specific microscopes imaging molecules at specific orientations $\bm{\Theta}$ [49], QFI provides a universal precision bound for any imaging system [46, 13]. While one cannot guarantee that a method can be devised to reach the QFI limit, comparing the CRB achieved by existing techniques to the quantum CRB (QCRB) [39] provides a useful context for evaluating performance.
@@
where $a$ incorporates the absorption cross-section and quantum yield of the molecule, wavelength of the excitation beam, etc. [39, 52] and $\bm{H}\in\mathbb{R}^{NL\times 15}$ represents the imaging system’s response to each fourth moment $q$, and models the impacts of excitation beam polarization and intensity and DSF engineering on the final images.
@@
Figure 4: Performance of using combined excitation modulation (ExM) and dipole-spread function (DSF) engineering. (a) Schematic of (i) 3 designed wavevectors $\bm{k}$ and (ii) their associated s- and p-polarized states for ExM. The three $\bm{k}$ vectors lie on a cone oriented along the $\mu_{z}$ direction with half-angle $\beta$, and the relative azimuthal angle between any pair is $120^{\circ}$. (b) Estimation precision for the (i) centroid orientation of and (ii) separation between a pair of dipoles using ExM-pixOL with various $\beta$ angles. Purple, $\beta=30^{\circ}$; orange, $\beta=45^{\circ}$; red, $\beta=\arccos{(1/\sqrt{3})}\approx 54.74^{\circ}$; blue, $\beta=60^{\circ}$. (c) Same as (b) for different DSFs and $\beta=\arccos{(1/\sqrt{3})}\approx 54.74^{\circ}$. Red, pixOL DSF [32]; green, polarized vortex DSF [30]; yellow, CHIDO [29]. The black line in (b) and (c) refers to the QCRB of fluorescence emission (ems, i.e., DSF engineering, Fig. 2). Precisions are calculated for a dipole pair with separation $\Delta=20^{\circ}$ and are averaged over $\phi_{c}\in[0^{\circ},360^{\circ}]$ with 100 photons detected and zero background.
@@
In summary, we analyzed the fundamental physical limits of measuring the orientations of a pair of overlapping dipole emitters using polarization-sensitive single-molecule imaging. Since the second moments of a pair of dipole emitters and a rotating molecule are identical, neither excitation modulation nor DSF engineering techniques can distinguish between the two scenarios without prior knowledge. However, by increasing the dimensionality of the measurement to the fourth-moment space via combined ExM and DSF engineering, pairs of dipoles produce camera images that are unique from those of a single wobbling molecule, thereby solving the degenerate measurements of existing techniques. We demonstrated that combined ExM-DSF engineering exhibits superior performance for measuring the centroid orientation of and separation between a pair of molecules. Our work lays the foundation for developing optimal techniques for resolving spatially overlapping dipole emitters and precisely measuring their rotational dynamics in the crowded environments typical of biological samples.
@@ Acknowledgements.
We thank Tingting Wu, Weiyan Zhou, and Suraj Deepak Khochare for helpful and insightful discussions. This work was supported by the National Institute of General Medical Sciences of the National Institutes of Health under Award Number R35GM124858.